OpenStudy (anonymous):

What is the solution to x=sqrt(-x+6)? I got 2

OpenStudy (haseeb96):

taking square on both sides x^2 = (-x+6) x^2+x-6=0 x^2+2x -3x -6=0 x(x+2) -3( x+2)=0 (x-3)(x+2)=0 either x-3=0 or x+2=0 x=3 or x =-2

OpenStudy (isaiah.feynman):